OpenStudy (rissaunnie):

FAN AND MEDAL What is the solution of the system defined by y = -x + 5 and 5x + 2y = 14?

OpenStudy (anonymous):

4/3, 11/3

OpenStudy (anonymous):

y=−x+5;5x+2y=14 y=−x+5for y: −x+5 5x+2y=14: 5x+2y=14 5x+2(−x+5)=14 3x+10=14 3x+10+−10=14+−10 3x=4 3x/3 = 4/3 x=4/3 Substitute 4/33 for x=−x+5: y=−x+5 y=−4/3+5 y=11/3